Calculate the particle velocity of air
Answer #1

a) Given p = vZ

p = 0.01 pascals , Z = 400rayls

v = p/Z

= 0.01/400

= 0.25 * 10^-4 m/s

b) Given V = 2(pi)fA

   To find amplitude A, A = v/(2(pi)f

   = (0.24 *10^-4)/ (2(pi)(1000))

   = 3.98 * 10^-9 N/m^2